# Optimal Mechanisms for Demand Response: An Indifference Set Approach

### v1

@@ 2 Optimal Demand Response
The optimal demand response problem involves implementing acceptable energy consumption profiles for each consumer such as to optimize a grid objective. If the grid has a cost function $\sigma(Q,\,Q_{0})$ for producing energy $Q\in\mathbb{R}^{d}$ given renewable production $Q_{0}$, then the optimal demand response problem requires finding a solution to
@@
Consider a consumer who only needs to add 50 kWh of charge to an electric vehicle by 7 am. Their indifference set is then $\mathbf{R}_{1}=\{q\geq 0:\sum_{j=1}^{7}q_{j}=50,\,q_{j}=0\text{ for all }j>7\}$.
@@
Consider a consumer who only wants operate an air conditioner to maintain the indoor temperature of their house within an acceptable range. Following Li et al. (2011), we model indoor temperature dynamics using the following linear formulation motivated by the heat equation,
@@
$\mathbf{R}_{3}=\left\{q\geq 0:T^{\mathsf{min}}\leq\sum_{i=1}^{j}\left(1-\alpha\right)^{i-j}\left(\beta q_{i}+\alpha T_{i}^{\mathsf{out}}\right)+(1-\alpha)^{i}\,T^{\mathsf{in}}_{0}\leq T^{\mathsf{max}}\text{ for all }1\leq j\leq 24\right\}.$
@@
Consider a consumer who wants to operate all 3 devices described above. Then their indifference set is $\mathbf{R}_{4}=\mathbf{R}_{1}\oplus\mathbf{R}_{2}\oplus\mathbf{R}_{3}$.
@@
At this point, our setting is still very abstract: For example, our definitions so far, it would, in principle, allow for the HEMS to give the grid operator direct control of all consumer devices, and let the grid solve the optimization problem (1). This approach would, of course, achieve optimal demand response, but would likely not be implementable in real-world systems. A more practical alternative is to decouple decision making by the grid and by the HEMS via a price mechanism: The HEMS optimizes consumption to minimize energy costs under (potentially time-varying) prices, and the grid chooses prices such as to induce desirable consumption profiles.
@@
resulting in an objective value $\widehat{\mathsf{R}}_{n}^{\mathsf{PM}}$. Note that, in the above definition, price-responsive HEMS are only responsive to variation in the price and not in the overall level of the price, i.e., $q_{i}(p)=q_{i}(\lambda\,p)$ for any $\lambda>0$. This reflects our modeling assumption that demand-response programs can only shift energy consumption within the indifference set that reflects the user’s pre-determined needs; they cannot induce consumers to revisit their needs. Similarly, optimal prices in (4) are scale invariant: If $p^{*}$ is optimal, then $\lambda\,p^{*}$ is also optimal for any $\lambda>0$. Thus, a successful demand response will only depend on the relative magnitude of prices across the day, and the overall price levels can be chosen to match revenue targets; see the discussion following Theorem 5 for more details.
We are now ready to state our main research questions: First, can a pricing mechanism as in (4) solve the optimal demand response problem (1), or will systems with pricing-based HEMS fall short of what could be achieved via direct device control by the grid? Second, if an optimal pricing mechanism exists, are there computationally efficient and practically applicable algorithms to learn the optimal prices? Overall, we will find that while price-based mechanisms may be suboptimal in finite systems, optimality gap of pricing as in (4) vanishes as the system size grows. Furthermore, optimal prices can be recovered via a simple first-order algorithm.
@@ 2.1 A Warm-up Illustration
In Figure 2, we plot the resulting $L_{n}(z)$ and $U_{n}(z)$ curves for both a realization of a small system with $n=10$ consumers and a realization of a larger system with $n=100$ consumers.
@@ 3 Mean-field Optimality of Dynamic Pricing
where $\mathsf{conv}$ denote the convex hull operation. [^4] Note that since $\mathsf{conv}(\mathfrak{C}_{d})$ is a closed subset of $\mathfrak{C}_{d}$ and $\mathsf{conv}$ is a continuous mapping with respect to the Hausdorff distance in a finite-dimensional space, $\mathsf{conv}(\mathbf{R})$ is measurable, and its expectation is well-defined. We are now ready to state our main assumptions.
@@
We note that the net-demand cost function $\rho(\cdot)$ can flexibly capture many types of costs. For example, cost functions that satisfy this assumption include $\rho(q)=\lVert q\rVert_{\infty}$ which measures peak demand; $\rho(q)=\smash{\sum_{j=1}^{d}|q_{j}|}$ which measures total demand; $\rho(p)=\smash{\lVert q\rVert_{\kappa}=(\sum_{j=1}^{d}q_{j}^{\kappa})^{1/\kappa}}$ for $\kappa>1$ which makes a compromise between these two; or $\rho(p)=\smash{\lVert q\rVert_{\kappa}+\lambda\sum_{j=1}^{d-1}|q_{j+1}-q_{j}|}$ with $\lambda>0$ which adds an extra penalty for fast ramp-ups or ramp-downs in demand. We now state some technical results that highlight key consequences of these assumptions. First, under our setting, set-expectation commutes with the user cost function in the following sense:
@@
Write the user-$i$ cost function from solving (4) as
@@
and, under Assumption 1, let $H(p)=h(p;\,\mathbb{E}\left[\mathbf{R}_{i}\right])$. Then, for any $p\in\mathbb{R}^{d}$ we also have $H(p)=\mathbb{E}\left[h(p;\,\mathbf{R}_{i})\right]$.
@@
(2) Whenever $\mathbf{R}_{i}$ has a non-atomic distribution, we have $\mathbb{E}\left[\mathsf{conv}(\mathbf{R}_{i})\right]=\mathbb{E}\left[\mathbf{R}_{i}\right]$ (Aumann, 1965; Artstein, 1974); and (3) the support function of a set and its convex hull are equal, i.e., $\delta(.;\,\mathbf{R})=\delta(.;\,\mathsf{conv}(\mathbf{R}))$. It then follows that
@@
$\begin{split}\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}&=\min\left\{\frac{1}{n}\sigma\left(\sum\limits_{i=1}^{n}q_{i},\,Q_{0}\right):q_{i}\in\mathbf{R}_{i}\text{ for all }1\leq i\leq n\right\}\\ &=\min\left\{\rho\left(\frac{1}{n}\sum\limits_{i=1}^{n}q_{i}-Q_{0}\right):q_{i}\in\mathbf{R}_{i}\text{ for all }1\leq i\leq n\right\}\\ &=\min\left\{\rho\left(q-Q_{0}/n\right):q\in\widehat{\mathbf{R}}_{n}\right\}.\end{split}$ (17)
@@
Furthermore, $\mathbb{E}\left[\mathbf{R}\right]$ is convex as argued in the proof of Lemma 2 and compact because $\mathbb{E}\left[\lVert\mathbf{R}\rVert\right]<\infty$, so by the minimax theorem
@@
The above result gives the scaled costs for both the direct and price-based mechanisms; it now remains to show that (15) and the minimizer of (16) over $p$ match. The function $\rho(\nabla H(p)-q_{0})$ is not necessarily convex in $p$ because it depends on $p$ through the differential of $H$; thus, classic duality theory for convex functions does not apply here. The following result nonetheless establishes optimality of the price-based mechanism under a differentiability condition.
@@
We assume that $h(z,\,\mathbf{R}_{i})$ as defined in (13) is almost surely differentiable at any point
@@
Theorem 4 above showed that, under considerable generality, pricing with prices of the form $p^{*}=\lambda\,z^{*}$ with $\lambda>0$ and $z^{*}$ as in (21) solves the demand-response problem. This abstract result highlights the promise of the pricing mechanism, but also raise some immediate practical questions such as
@@ 4 A First-order Algorithm for Optimal Pricing
Under Assumptions 1 and 2, let $\hat{z}_{n}$ be any sequence of prices satisfying (27) for $n$ number of consumers. Then
@@
The next question is then how to efficiently maximize the function $\widehat{H}(z)$. At face value, the problem (27) looks like a zero-th order optimization problem: We seek to maximize a random function by taking draws at different parameter values. In general, zero-th order problems are solvable but not particularly well-behaved. In the best case scenario where we can noiselessly query the same unit multiple times, the error bound of zero-th order optimization are a factor $\sqrt{d}$ worse than what could typically be achieved via 1-st order methods with access to gradients (Duchi et al., 2015). [^8] These methods proceed by effectively taking numerical gradients by evaluating $h(z,\,\mathbf{R}_{i}),\,h(z^{\prime},\,\mathbf{R}_{i})$ for nearby values of $z,\,z^{\prime}$. In addition, the situation gets much worse if there’s any noise when the grid queries each HEMS to get $h(z,\,\mathbf{R}_{i})$ (or if it’s only possible to query each HEMS once); in this case, zero-th order optimization falls short of 1-st order behavior even in terms of its rate of convergence (Shamir, 2013).
@@
  • Form a gradient $g_{k}=\bar{q}_{k}-q_{0}$, and use it to update $z_{k}$.
There is a huge number of available 1-st order algorithms that could be used here. For example, one could use standard gradient descent algorithms that repeatedly query all consumers (i.e., set $\mathcal{N}_{k}=\left\{1,\,\ldots,\,n\right\}$), or stochastic gradient descent algorithms that learn by sequentially going through the sample and only querying each customer once (i.e., with $\mathcal{N}_{k}=\left\{k\right\}$).
@@ 5 Case Study: Pre-cooling in Phoenix, Arizona
In this section, we examine a sample grid of households equipped with Home Energy Management Systems (HEMS) on a summer day. To simplify the analysis, we suppose that household electricity usage is solely for cooling. In this setting, homeowners input a desired range of indoor temperatures into their energy management systems, and then HEMS calculates a power trajectory based on the outside temperature, the house’s thermal properties, user’s preferences, and the electricity price.
For the outside temperature, we use the average hourly temperature data [^10] https://www.visualcrossing.com/weather/weather-data-services of Phoenix, Arizona in July 2023. We suppose that summer days in Phoenix are relatively consistent, using data from $n=31$ days to get the sample mean $\hat{\mu}_{d}^{\mathsf{temp}}$ and the sample covariance matrix $\hat{\Sigma}_{d}^{\mathsf{temp}}$ for $d=24$ hours. In the next step, we follow the indoor temperature dynamics given as (2) in Example 3. In our model, we assume that the outside temperature is the same for all households, and that each house has a fixed preference to maintain its inside temperature within the interval [20, 25] degrees, with an initial temperature of $T^{\mathsf{in}}_{0}=24$ degrees. Following the temperature dynamics (2), we suppose heterogeneity across users for thermal parameters $\alpha_{i}$ and $\beta_{i}$, and we let
@@
We next use Arizona’s hourly renewable energy data [^11] https://www.eia.gov/opendata/browser/electricity/rto for solar in July 2023 to obtain the sample mean $\hat{\mu}_{d}^{\mathsf{renewable}}$ and the sample covariance matrix $\hat{\Sigma}_{d}^{\mathsf{renewable}}$ for available renewable.
We consider the following grid cost function parametrized by $s\geq 1$:
@@
For obtained prices $p^{*}_{s}$, we follow the HEMS response function (3) to obtain users consumption profiles. We first compute the total demand and inside temperatures in response to the price signal $p^{*}_{s}$ for $s=1,2,4,$ and +$\infty$+ ∞. The results are shown in Figures 3(c) and 3(d) respectively for net demand and inside temperatures. Under dynamic pricing, the average indoor temperatures exhibit a pre-cooling phase before reaching the high-demand periods, resulting in a different demand curve. In addition, a detailed view of temperature profiles of all users is exhibited in Figure 3(e). For $\sigma_{1},\sigma_{2},\sigma_{4}$, and $\sigma_{\infty}$ the total grid cost function under flat-rate pricing is respectively 2147.56, 619.29, 347.21, and 236.12, whereas when the optimal dynamic prices $p^{*}_{s}$ are deployed, the obtained total grid cost values are respectively 2041.93, 515.96, 250.88, and 153.52. This implies approximately 4.9%, 16.7%, 27.7%, and 35.0% cost savings, respectively.
@@
(b) Average voltage variations of different pricing signals
@@
(d) Inside temperatures and averages for $p^{*}_{s}$
(e) Detailed inside temperatures and averages
Figure 3: Results of numerical experiments evaluating the performance of dynamic pricing. Figure 3(a) shows the optimal price signal $p^{*}_{s}$ obtained using the Frank-Wolfe Algorithm for various values of the cost function $\sigma_{s}(.)$. Figure 3(b) shows voltage variations for $d$ periods across different pricing mechanisms. Figures 3(c) and 3(d) display the demand curves and inside temperature profiles under $p^{*}_{s}$ and flat-rate pricing, respectively. Figure 3(e) provides a detailed (zoomed-in) view of Figure 3(d). Overall, deploying the dynamic pricing $p^{*}_{s}$ instead of flat-rate pricing results in cost savings of approximately 4.9%, 16.7%, 27.7%, and 35.0% for the grid cost functions $\sigma_{1}$, $\sigma_{2}$, $\sigma_{4}$, and $\sigma_{\infty}$, respectively.
@@
In addition, we keep the physical thermal properties of the houses, $\alpha$ and $\beta$, unchanged. Each day, we conduct four experiments, observing users’ HEMS responses to four pricing signals: $p^{*}_{1}$, $p^{*}_{2}$, $p^{*}_{4}$, and $p^{*}_{\infty}$. Our goal is to assess the effectiveness of $p^{*}_{s}$ in reducing the grid cost function $\sigma_{s}(\cdot)$ compared to flat-rate pricing. The exact percentage reductions in cost values compared to flat-rate pricing, calculated across all days and for various grid cost functions, are presented in Figure 4(a). In addition, a summary of these results is shown in Figure 4(b), where violin plots illustrate the outcomes for four values of $s=1,2,4,$ and +$\infty$+ ∞ over 45 days. It can be observed that even with fluctuations in outside temperature, the previously computed prices are effective in reducing the grid cost function. This effect becomes more pronounced as $s$ increases, although with greater variation. We further demonstrate the total demand curves under $p^{*}_{s}$ and flat-rate pricing for all 45 days in Figure 4(c). It is evident that the duck-shaped net-demand curve under flat-rate pricing is effectively flattened under dynamic pricing.
Finally, we compute the temperature profiles under $p^{*}_{s}$ for various values of $s$ and for flat-rate pricing across all 45 days, with the results displayed in Figures 5(a). To better illustrate the variation among different days, we separately plot the inside temperature dynamics in Figure 5(b).
@@
(b) Violin plots of percentage decline in grid cost $\sigma_{s}$ when using $p^{*}_{s}$ compared to flat-rate pricing over 45 days

### v2

@@ 2 Optimal Demand Response
To design effective demand response mechanisms, we first establish a framework that captures consumer flexibility. We introduce the concept of indifference sets, which define the range of consumption profiles consumers find acceptable, and formulate the optimization problem faced by both the grid and consumers.
@@
The optimal demand response problem involves implementing acceptable energy consumption profiles for each consumer such as to optimize a grid objective. If the grid has a cost function $\sigma(Q,\,Q_{0})$ for producing energy $Q\in\mathbb{R}^{d}$ given renewable production $Q_{0}$, then the optimal demand response problem requires finding a solution to the Direct Method (DM) risk value
@@
Consider a consumer who plugs in an electric vehicle at midnight and only needs to add 50 kWh of charge by 7 am. Their indifference set is then $\mathbf{R}_{1}=\{q\geq 0:\sum_{j=1}^{7}q_{j}=50,\,q_{j}=0\text{ for all }j>7\}$.
@@
Consider a consumer who only wants to operate an air conditioner to maintain the indoor temperature of their house within an acceptable range. Following Li et al. (2011), we model indoor temperature dynamics using the following linear formulation motivated by the heat equation,
@@
$\mathbf{R}_{3}=\left\{q\geq 0:T^{\mathsf{min}}\leq\sum_{i=1}^{j}\left(1-\alpha\right)^{j-i}\left(\beta q_{i}+\alpha T_{i}^{\mathsf{out}}\right)+(1-\alpha)^{j}\,T^{\mathsf{in}}_{0}\leq T^{\mathsf{max}}\text{ for all }1\leq j\leq 24\right\}.$
@@
Consider a consumer who wants to operate all 3 devices described above. Then their indifference set is $\mathbf{R}_{4}=\mathbf{R}_{1}\oplus\mathbf{R}_{2}\oplus\mathbf{R}_{3}$. [^13] The Minkowski sum (vector sum) of two indifference sets defined as $\mathbf{R}_{1}\oplus\mathbf{R}_{2}=\{a+b:a\in\mathbf{R}_{1},b\in\mathbf{R}_{2}\}$.
@@
At this stage, our framework remains abstract: For example, our definitions so far, it would, in principle, allow for the HEMS to give the grid operator direct control of all consumer devices, and let the grid solve the optimization problem (1). This approach would, of course, achieve optimal demand response, but would likely not be implementable in real-world systems. A more practical alternative is to decouple decision making by the grid and by the HEMS via a price mechanism: The HEMS optimizes consumption to minimize energy costs under (potentially time-varying) prices, and the grid chooses prices such as to induce desirable consumption profiles.
@@
resulting in an objective value $\widehat{\mathsf{R}}_{n}^{\mathsf{PM}}$ as the optimal Price-based Method (PM) risk value. We are now ready to state our main research questions: First, can a pricing mechanism as in (4) solve the optimal demand response problem (1), or will systems with pricing-based HEMS fall short of what could be achieved via direct device control by the grid? Second, if an optimal pricing mechanism exists, are there computationally efficient and practically applicable algorithms to learn the optimal prices? Overall, we will find that while price-based mechanisms may be suboptimal in finite systems, optimality gap of pricing as in (4) vanishes as the system size grows. Furthermore, optimal prices can be recovered via a simple first-order algorithm.
Remark 1.
In our model, consumer indifference sets remain exogenous to prices, i.e., consumers respond to prices by optimizing their consumption with $\mathbf{R}_{i}$ as in (3), but the $\mathbf{R}_{i}$ themselves remain fixed. This modeling choice reflects our underlying research question: How can we design mechanisms that leverage existing consumer flexibility for demand response (and how does a price-based mechanism compare with direct control)? Once the $\mathbf{R}_{i}$ are fixed and coded into HEMS, optimal demand response essentially becomes an optimization problem, and our challenge is to find (or approximate) the optimum in a feasible and pragmatic way (i.e., in a way that respects natural communication and privacy desiderata). Reasoning about how the $\mathbf{R}_{i}$ themselves may change in the long run in response to new pricing strategies is an interesting question (see Section 6 for further discussion); however, doing so would likely not be automatable using HEMS (and would instead require a deeper understanding of how consumers update preferences over time), and so is outside the scope of the present analysis.
Remark 2.
As a corollary to the fact that the indifference sets $\mathbf{R}_{i}$ are exogenous with respect to prices we observe that, in our model, price-responsive HEMS are only responsive to variation in the shape of the price curve and not in the overall level of the price, i.e., $q_{i}(p)=q_{i}(\lambda p)$ for any $\lambda>0$. Similarly, optimal prices in (4) are scale invariant: If $p^{*}$ is optimal, then $\lambda\,p^{*}$ is also optimal for any $\lambda>0$. Thus, in our model, demand response is purely about cross-time substitution of energy use, and successful price-based demand response involves understanding how the the relative magnitudes of the price signal over time will shape demand.
Remark 3.
Because behavior is scale-invariant to prices, the overall price level remains a free parameter in our model—and can be set by the regulator and/or the utilities. In other words, once our approach outputs a price curve $p^{*}$, the regulator can rescale it to $\lambda\,p^{*}$ for any $\lambda>0$ in order to control some aggregate metrics of energy costs or utility revenue; see the discussion following Theorem 5 for more details.
@@ 2.1 A Warm-up Illustration
In Figure 2, we plot the resulting $n^{-1}L_{n}(z)$ and $n^{-1}U_{n}(z)$ curves for both a realization of a small system with $n=10$ consumers and a realization of a larger system with $n=100$ consumers.
@@ 3 Mean-field Optimality of Dynamic Pricing
With the indifference set framework in place, we examine whether price-based mechanisms can achieve optimal demand response. We show that while pricing is generally suboptimal in finite markets, it converges to optimality in a mean-field setting as the number of consumers grows.
@@
where $\mathsf{conv}$ denote the convex hull operation. [^15] Note that since $\mathsf{conv}(\mathfrak{C}_{d})$ is a closed subset of $\mathfrak{C}_{d}$ and $\mathsf{conv}$ is a continuous mapping with respect to the Hausdorff distance in a finite-dimensional space, $\mathsf{conv}(\mathbf{R})$ is measurable, and its expectation is well-defined. We are now ready to state our main assumptions. Our first assumption is a standard technical assumption, and implies useful concentration properties for random sets.
@@
Next, we introduce an assumption the grid risk functions. In real-world applications, practitioners often aim to regulate peak demand, total consumption, or sharp transitions between periods. We find that a class of order-one positively homogeneous grid risk functions is sufficiently powerful to capture these behaviors and effectively shape the net-demand curve.
@@
We now state some technical results that highlight key consequences of these assumptions. First, under our setting, set-expectation commutes with the user cost function in the following sense:
@@
Under Assumption 1, write the user-$i$ cost function from solving (4) as
@@
and let $H(p)=h(p;\,\mathbb{E}\left[\mathbf{R}_{i}\right])$. Then, for any $p\in\mathbb{R}^{d}$ we also have $H(p)=\mathbb{E}\left[h(p;\,\mathbf{R}_{i})\right]$.
@@
(2) Whenever $\mathbf{R}_{i}$ has a non-atomic distribution, we have $\mathbb{E}\left[\mathsf{conv}(\mathbf{R}_{i})\right]=\mathbb{E}\left[\mathbf{R}_{i}\right]$ (Aumann, 1965); and (3) the support function of a set and its convex hull are equal, i.e., $\delta(.;\,\mathbf{R})=\delta(.;\,\mathsf{conv}(\mathbf{R}))$. It then follows that
@@
$\begin{split}\frac{1}{n}\widehat{\mathsf{R}}_{n}^{\mathsf{DM}}&=\min\left\{\frac{1}{n}\sigma\left(\sum\limits_{i=1}^{n}q_{i},\,Q_{0}\right):q_{i}\in\mathbf{R}_{i}\text{ for all }1\leq i\leq n\right\}\\ &=\min\left\{\rho\left(\frac{1}{n}\sum\limits_{i=1}^{n}q_{i}-Q_{0}/n\right):q_{i}\in\mathbf{R}_{i}\text{ for all }1\leq i\leq n\right\}\\ &=\min\left\{\rho\left(q-Q_{0}/n\right):q\in\widehat{\mathbf{R}}_{n}\right\}.\end{split}$ (17)
@@
Furthermore, because we are working in a non-atomic probability space and $\mathbb{E}\left[\lVert\mathbf{R}\rVert\right]<\infty$, the set $\mathbb{E}\left[\mathbf{R}\right]$ must be convex and compact (Aumann, 1965), and so by the minimax theorem
@@
The above result gives the scaled costs for both the direct and price-based mechanisms; it now remains to show that (15) and the minimizer of (16) over $p$ match. The function $\rho(\nabla H(p)-q_{0})$ is not necessarily convex in $p$ because it depends on $p$ through the differential of $H$; thus, classic duality theory for convex functions does not apply here. The following result nonetheless establishes optimality of the price-based mechanism under one further assumption: We require sufficient stochasticity in the $\mathbf{R}_{i}$ for the mean-field cost function ${\mathbb{E}}[h(p;\mathbf{R})]$ to be smooth and differentiable at its maximizer(s). We note that this phenomenon (i.e., differentiability at the maximizer) is generally not true for the individual cost functions themselves. [^17] We emphasize that this randomness pertains to the distribution of consumption profiles $\mathbf{R}_{i}$ rather than consumer preferences. For example, in our Phoenix experiment discussed in Section 5, all consumers have similar preferences—yet heterogeneity in house characteristics (e.g., differences in home insulation properties, cooling system performance, etc.) provide sufficient randomness to justify Assumption 3.
@@
The cost $h(z,\,\mathbf{R}_{i})$ as in (13) is almost surely differentiable at any solution to
@@
Theorem 4 above showed that, under considerable generality, pricing with prices of the form $p^{*}=\lambda\,z^{*}$ with $\lambda>0$ and $z^{*}$ as in (21) solves the demand-response problem. This abstract result highlights the promise of the pricing mechanism, but also raise some immediate practical questions such as:
@@ 4 A First-order Algorithm for Optimal Pricing
Given the theoretical foundation for pricing mechanisms, we develop a practical algorithm for computing optimal dynamic prices. Our approach relies on first-order optimization methods and requires only minimal information from HEMS, making it both efficient and scalable.
@@
Under Assumptions 1 and 2, let $\hat{z}_{n}$ be any sequence of prices satisfying (27) for $n$ consumers. Then
@@
The next question is then how to efficiently maximize the function $\widehat{H}(z)$. At face value, the problem (27) looks like a zeroth-order optimization problem: We seek to maximize a random function by taking draws at different parameter values. In general, zeroth-order problems are solvable but not particularly well-behaved. In the best case scenario where we can noiselessly query the same unit multiple times, the error bound of zeroth-order optimization are a factor $\sqrt{d}$ worse than what could typically be achieved via first-order methods with access to gradients (Duchi et al., 2015). [^20] These methods proceed by effectively taking numerical gradients by evaluating $h(z,\,\mathbf{R}_{i}),\,h(z^{\prime},\,\mathbf{R}_{i})$ for nearby values of $z,\,z^{\prime}$. In addition, the situation gets much worse if there’s any noise when the grid queries each HEMS to get $h(z,\,\mathbf{R}_{i})$ (or if it’s only possible to query each HEMS once); in this case, zeroth-order optimization falls short of first-order behavior even in terms of its rate of convergence (Shamir, 2013).
@@
  • Form a gradient $g_{k}=\bar{q}_{k}-q_{0}$, and use it to update $z_{k}$ and obtain a new $z_{k+1}\in\mathcal{P}$.
There is an extensive literature on first-order algorithms that can be used to implement this strategy (Bertsekas, 1999; Parikh et al., 2014). For example, one could use standard projected or constrained gradient descent algorithms that repeatedly query all consumers (i.e., set $\mathcal{N}_{k}=\left\{1,\,\ldots,\,n\right\}$), or stochastic alternatives that learn by sequentially going through the sample and only querying each customer once (i.e., with $\mathcal{N}_{k}=\left\{k\right\}$).
@@ 5 Case Study: Pre-cooling in Phoenix, Arizona
In this section, we examine a sample grid of households equipped with Home Energy Management Systems (HEMS) on a summer day from Phoenix, Arizona, a particularly relevant testbed due to its extreme heat and abundant solar generation. To simplify the analysis, we only considered household electricity usage for cooling. In this setting, homeowners input a desired range of indoor temperatures into their energy management systems, and then HEMS calculates a power trajectory based on the outside temperature, the house’s thermal properties, user’s preferences, and the electricity price.
For the outside temperature, we use the average hourly temperature data [^22] https://www.visualcrossing.com/weather/weather-data-services of Phoenix, Arizona in July 2023. We suppose that summer days in Phoenix are relatively consistent, using data from $n=31$ days to get the sample mean $\hat{\mu}_{d}^{\mathsf{temp}}$ and the sample covariance matrix $\hat{\Sigma}_{d}^{\mathsf{temp}}$ for $d=24$ hours. In the next step, we follow the indoor temperature dynamics as given in Example 3, i.e., $T^{\mathsf{in}}_{it}=T^{\mathsf{in}}_{i(t-1)}+\alpha_{i}(T^{\mathsf{out}}_{it}-T^{\mathsf{in}}_{i(t-1)})+\beta_{i}q_{it}$, where $\alpha_{i}\in\mathbb{R}$ characterizes the $i$-th house’s thermal insulation, and $\beta_{i}<0$ quantifies its cooling system’s effectiveness. In our model, we assume that the outside temperature is the same for all households, and that each house has a fixed preference to maintain its inside temperature within the interval [20, 25] degrees, with an initial temperature of $T^{\mathsf{in}}_{0}=24$ degrees. Following the temperature dynamics (2), we suppose heterogeneity across users for thermal parameters $\alpha_{i}$ and $\beta_{i}$, and let [^23] These choices for $\alpha_{i}$ and $\beta_{i}$ are made to obtain proportionally meaningful energy levels for the consumption of HVAC systems over one hour. We also intentionally choose small $\alpha_{i}$ values (indicating good insulation properties) to better accommodate pre-cooling.
@@
We use Arizona’s hourly renewable energy data [^24] https://www.eia.gov/opendata/browser/electricity/rto for solar in July 2023 to obtain the sample mean $\hat{\mu}_{d}^{\mathsf{renewable}}$ and the sample covariance matrix $\hat{\Sigma}_{d}^{\mathsf{renewable}}$ for available renewable. We consider the following grid cost function parametrized by $s\geq 1$:
@@
For obtained prices $p^{*}_{s}$, we follow the HEMS response function (3) to obtain users consumption profiles. We first compute the total demand and inside temperatures in response to the price signal $p^{*}_{s}$ for $s=1,2,4,$ and +$\infty$+ ∞. The results are shown in Figures 3(c) and 3(d) respectively for net demand and inside temperatures. Under dynamic pricing, the average indoor temperatures exhibit a pre-cooling phase before reaching the high-demand periods, resulting in a different demand curve. In addition, a detailed view of the temperature profiles of all consumers under different price signals is shown in Figure 4, where, in each sub-figure, thinner curves represent the indoor temperature dynamics of individual consumers throughout a full day. It can be observed that the optimal price for the grid risk function $\sigma_{s}(.)$ incentivizes a larger portion of consumers, and more strongly, to pre-cool their houses for larger values of $s$. The major pre-cooling occurs during midday hours (10–15), when abundant renewable energy is available, which helps mitigate the sharp ramp-up during the late evening around 20:00. For $\sigma_{1},\sigma_{2},\sigma_{4}$, and $\sigma_{\infty}$ the total grid cost function under flat-rate pricing is respectively 2147.56, 619.29, 347.21, and 236.12, whereas when the optimal dynamic prices $p^{*}_{s}$ are deployed, the obtained total grid cost values are respectively 2041.93, 515.96, 250.88, and 153.52. This implies approximately 4.9%, 16.7%, 27.7%, and 35.0% cost savings, respectively.
@@
(b) Average voltage variations of different prices
@@
(d) Inside average temperatures for $p^{*}_{s}$
Figure 3: Results of numerical experiments evaluating the performance of dynamic pricing. Figure 3(a) shows the optimal price signal $p^{*}_{s}$ obtained for a variety of grid risk function $\sigma_{s}(.)$. Figure 3(b) shows voltage variations for $d$ periods across different pricing mechanisms. Figures 3(c) and 3(d) display the demand curves and inside temperature profiles under $p^{*}_{s}$ and flat-rate pricing, respectively. Overall, deploying the dynamic pricing $p^{*}_{s}$ instead of flat-rate pricing results in cost savings of approximately 4.9%, 16.7%, 27.7%, and 35.0% for the grid cost functions $\sigma_{1}$, $\sigma_{2}$, $\sigma_{4}$, and $\sigma_{\infty}$, respectively.
(a) Flat Rate
(b) $p^{*}_{1}$
(c) $p^{*}_{2}$
(d) $p^{*}_{4}$
(e) $p^{*}_{\infty}$
Figure 4: Comparison of all consumers’ temperature profiles under different pricing strategies, $p^{*}_{s}$—the optimal price for the grid risk function $\sigma_{s}(.)$. This figure provides a detailed (zoomed-in) view of Figure 3(d). Each thin curve represents the indoor temperature dynamics of an individual consumer.
@@
We keep the physical thermal properties of the houses, $\alpha$ and $\beta$, unchanged. Each day, we conduct four experiments, observing users’ HEMS responses to four pricing signals: $p^{*}_{1}$, $p^{*}_{2}$, $p^{*}_{4}$, and $p^{*}_{\infty}$. Our goal is to assess the effectiveness of $p^{*}_{s}$ in reducing the grid cost function $\sigma_{s}(\cdot)$ compared to flat-rate pricing. The exact percentage reductions in cost values compared to flat-rate pricing, calculated across all days and for various grid cost functions, are presented in Figure 5(a). In addition, a summary of these results is shown in Figure 5(b), where violin plots illustrate the outcomes for four values of $s=1,2,4,$ and +$\infty$+ ∞ over 45 days. It can be observed that even with fluctuations in outside temperature, the previously computed prices are effective in reducing the grid cost function. This effect becomes more pronounced as $s$ increases, although with greater variation. We further demonstrate the total demand curves under $p^{*}_{s}$ and flat-rate pricing for all 45 days in Figure 5(c). It is evident that the duck-shaped net-demand curve under flat-rate pricing is effectively flattened under dynamic pricing.
Finally, we compute the temperature profiles under $p^{*}_{s}$ for various values of $s$ and for flat-rate pricing across all 45 days, with the results displayed in Figures 6(a). To better illustrate the variation among different days, we separately plot the inside temperature dynamics in Figure 6(b). We see that our proposed algorithm in fact lowers indoor temperatures during the day when solar power is available—as one would conceptually expect with pre-cooling.
@@
(b) Violin plots of percentage decline in grid risk when using $p^{*}_{s}$ compared to flat-rate pricing